# Defect recombination origin of low energy excess in semiconductor detectors

### v1

@@ IV Avalanche mechanism of low-temperature annealing
The observation described in subsection III.3 of annealing occuring at cryogenic temperatures may seem surprising, as the temperature of 100 K corresponds to only roughly 0.01 eV/atom, much lower than the chemical bond energy of roughly 2 eV/bond. It is also not consistent with the traditional picture of radiation damage annealing being described to be caused by point defect migration with well-defined migration barriers.
@@
However, the result can qualitatively be understood as follows. The complex disordered damage pockets (cf. Fig. 2) will inadvertently have metastable bonds with a much lower energy barrier to recombination. Furthermore, once a metastable bond recombines towards the ground state energy, it can induce a localized recombination “chain reaction”: one bond starts annealing, then the released energy heats up the sample locally which can induce additional bond recombination. To illustrate this argument with a rough calculation, even a low energy release of 2 eV of one recombining bond corresponds to a transient heating of the nearest ten atoms by about 1000 K. To quantify this deduction, we now analyze in atomic-level detail what kind of mechanisms can lead to the energy releases of several eV’s even when the temperature is so low that thermal activation cannot possibly lead to crossing an energy barrier in the eV range.
@@
To do the analysis, some of the the 2 ps intervals where energy release events took place were picked out. Annealing simulations were restarted at the beginning of these intervals, and atomic information was recorded every 10 fs. The atomic configuration extracted every 10 fs became the input for quenching. Quenching was performed in the same way as before (Section II) to find out energy minima along the reaction path. Finally, transition state search by NEB method was carried out between neighboring energy minima to find out the energy barriers.
@@
Fig. 10 and Fig. 11 shows two energy release events whose energy release are larger than 8 eV, taking place at 0.05 K (QTB thermostat). In Fig. 10 and Fig. 11, (c)(d)(e)(f) shows the atomic motion. Both of the events go through multiple steps of energy drops, instead of releasing all energy at once. From Fig. 10 (b) and Fig. 11 (b) we see that the energy barriers of the fundamental steps are lower than 0.1 eV, making it possible for the energy releases to take place at very low temperatures. Moreover, both events involve the migration of several atoms, not just one or two of them.
@@
Figure 10: An example of the energy barriers in an individual annealing event. In this case the total energy release was 10.24 eV. In (c)(d)(e)(f), the color coding is atomic displacement.
@@
The results show that at cryogenic temperature, all energy barriers for annealing are very small, and most, if not all of them, are below 0.1 eV. If a metastable configuration overcomes a small energy barrier at first, additional atom rearrangement nearby can be triggered, and much more energy can be released, which can be viewed as an “avalanche” effect. In other words, the effect is a variety of self-organized critical phenomena, well known e.g to explain dynamics in sandpiles [65, 66].
@@
Since the primary radiation damage event will lead to a multitude of complex damage configurations, there is no minimum temperature at which annealing can occur, as there will be barriers of all height present. Therefore, energy release avalanches can be triggered at any temperature.
V Discussion and conclusion
The main result of this paper is that recombination of damage produced by radioactive events can cause energy release on very long time scales after the damage production, even at cryogenic temperatures. The simulations in this work were done on silicon, but considering that it is known that other covalently bonded semiconductor materials such as Ge and GaAs produce damage in very similar (and even larger) damage pockets as Si [23, 24, 67, 68], it is very likely that very similar behavior would occur in also other covalently bonded semiconductors.
@@ V Discussion and conclusion
Many ionic materials also undergo amorphization [69, 70, 71], and hence a similar mechanisms might be active in at least some of them. Elemental metals do not amorphize [72], however, metastable damage pockets can be formed in them in the form of ordered damage clusters (e.g. C15 structure) and dislocation loops [73, 74, 75, 76]. However, separate simulation work with appropriate interatomic interaction models for ionic and metallic materials would be needed to deduce whether energy release from such structures would have a similar energy dependence.
@@
The effect described in this paper has a time dependence on the ns to tens of nanoseconds time scale. Most experiments observing the low-energy excess have not reported such time dependence. This is most likely because their time resolution is large (of the order of ms). Hence if there is sufficient amounts of radioactive impurities or any other damage sources in the sample that numerous decays occur in the detector within the time resolution, the average signal on the long time scales would appear constant, as the detector effectively sums up multiple events. On the other hand, we do note that the predicted exponential time dependence of the energy release events on time scales of tens of nanoseconds at room temperature might be experimentally testable with time-resolved pulsed ion beam experiments [77, 78] and sufficiently fast detectors.
Moreover, as we have shown earlier, a dark matter signal could have a distinct daily variation [3, 79, 80], which the excess signals from radioactive decays would not have.
@@
In conclusion, using molecular dynamics simulations of radiation events and their annealing in silicon, we have shown that even at cryogenic temperatures, energy release events of up to 10’s of electron volts can be caused by random annealing of complex disordered atom pockets. We show that the recombination can occur by avalanche-like events, where a metastable defect configuration first crosses an energy barrier of about 0.1 eV. Crossing this kind of small barrier can then trigger a rapid additional bond recombinations, that lead to partial recrystallization and energy releases exceeding 10 eV. This explains why large energy release events can occur even down to cryogenic temperatures. These events occur randomly up to macrosopic time scales after the initial picosecond time scale radiation event. Hence, in a sensitive detector they could be caused by annealing of damage produced e.g. by decay of radioactive impurities either in the detector itself or in materials surrounding it. These energy release events have an exponential energy dependence that agrees very well with the experimentally observed low-energy excess energy release in semiconductor detectors.

### v2

@@ IV Avalanche mechanism of low-temperature annealing
The observation described in subsection III.3 and subsection III.4 of annealing occuring at cryogenic temperatures may seem surprising, because the temperature of 100 K corresponds to only roughly 0.01 eV/atom, much lower than the chemical bond energy of roughly 2 eV/bond. It is also not consistent with the traditional picture of radiation damage annealing, which is usually described to be caused by point defect migration with well-defined migration barriers.
@@
However, the result can qualitatively be understood as follows. The complex disordered damage pockets (cf. Figure 1) can inadvertently have metastable bonds with a much lower energy barrier to annealing. Furthermore, once a metastable bond recombines towards the ground state energy, it can induce a localized annealing “chain reaction”: one bond starts annealing, then the released energy heats up the sample locally which can induce additional bond rearrangements. To illustrate this argument with a rough calculation, even a low energy release of 2 eV of one recombining bond corresponds to a transient heating of the nearest ten atoms by about 1000 K. To quantify this deduction, we now analyze in atomic-level detail what kind of mechanisms can lead to the energy releases of several eV’s even when the temperature is so low that thermal activation cannot possibly lead to crossing an energy barrier in the eV range.
@@
To do the analysis, some of the the 2 ps intervals where energy release events took place were picked out. Annealing simulations were restarted at the beginning of these intervals, and atomic information was recorded every 10 fs. The atomic configuration extracted every 10 fs became the input for quenching. Quenching was performed in the same way as before (Section II) to find out energy minima along the reaction path. Finally, transition state search by climbing image nudged elastic band (CI-NEB) method was carried out between neighboring energy minima to find out the energy barriers.
Figure 10 and Figure 11 shows two energy release events whose energy release are larger than 8 eV, taking place at 0.05 K (QTB thermostat). Figure 10 (a) and Figure 11 (a) shows the potential energy and the number of atoms in the amorphous phase as functions of time. To get these figures, all atomic configurations have gone through quenching. The number of atoms in the amorphous phase is obtained by DXA analysis in Ovito [84]). Figure 10 (b) and Figure 11 (b) show the energy barriers of the fundamental steps in the two events, which are obtained by climbing image nudged elastic band (CI-NEB) method. Figure 10 (c)(d)(e)(f) and Figure 11 (c)(d)(e)(f) show the motion of atoms in these two events by color coding of atomic displacement (The zero point of atomic displacement is the atomic configuration before the energy release event). In each figure, the states in (c)(d)(e)(f) take place chronologically, from state 1 in (c) to state 4 in (f). These four states correspond to the four energy minima along the reaction path of the energy release events.
@@
From Figure 10 (a) and Figure 11 (a), one can see that instead of releasing energy all at once, both of the energy release events involve three fundamental steps and four energy minima along the reaction paths. One can observe that the energy releases in the events are often accompanied by the reduction of the number of atoms in the amorphous phase, namely, the crystallization of amorphous pockets. From Figure 10 (b) and Figure 11 (b) we see that most, if not all the energy barriers of the fundamental steps are lower than 0.1 eV, making it possible for the energy releases to take place at very low temperatures. Moreover, from Figure 10 (c)(d)(e)(f) and Figure 11 (c)(d)(e)(f), both events involve the migration of several atoms, not just one or two of them. This means that the evolution of defects in these energy release events is not the migration of point defects, but the rearrangement of many atoms.
@@
Figure 10: An example of the energy barriers in an individual annealing event. In this case the total energy release was 10.24 eV. In (c)(d)(e)(f), the color coding is the atomic displacement.
@@
The results show that at cryogenic temperature, all energy barriers for annealing are very small, and most, if not all of them, are below 0.1 eV. If a metastable configuration overcomes a small energy barrier at first, additional atom rearrangement nearby can be triggered due to the heat released in the first step (overcoming the initial small barrier) in the defect evolution process, and much more energy can be released, which can be viewed as an “avalanche” effect, just like a terrible avalanche on a snow mountain being triggered by tiny perturbations somewhere on the hill. In other words, the effect is a variety of self-organized critical phenomena, which are well known e.g to explain dynamics in sandpiles [85, 86].
@@
Since the primary radiation damage event will lead to a multitude of complex damage configurations, there is no minimum temperature at which annealing can occur, since there will be barriers of all height present. Therefore, energy release avalanches can be triggered at any temperature. Moreover, when compared with the results of Béland et al. and Jay et al. [35, 36, 37] showing relaxation events up to second timescales with the k-ART approach, one can conclude that annealing events may occur up to fully macroscopic times.
V Analysis of location of annealing events
To analyze where the energy is released in annealing, six annealing cases at 0.05 K were selected, and in total 35 energy release events were collected from these six cases. We visualized all of these events, adding color coding to the atom displacements, to find the atoms that moved the most in energy release events. The reference zero for atom displacements is the atom positions at the beginning of the 2 ps annealing interval where the energy release event takes place. We also used defect meshes from DXA analysis in Ovito [84] to mark the boundaries of amorphous pockets in silicon.
From the visualizations, we discovered that the energy release events took place in the amorphous pockets in damaged silicon. Some of them took place at large amorphous pockets with diameters larger than 10 Å, while others took place at quite small amorphous zones. Some of the energy release events took place at the boundaries of the amorphous pockets, which is the amorphous-crystalline interface. The other energy release events took place inside the amorphous pockets. Both of these situations can be viewed as rearrangements of atoms in the amorphous pockets.
A few examples are shown in the following figures. In these figures, atoms are colored by their displacements, and white meshes (defect meshes) mark the boundaries of amorphous pockets. Figure 12 (a) shows an energy release event that takes place at the boundary of a large amorphous pocket, where the proximity of the atom movements to the pocket boundary is clearly shown by the colors of the atoms and the white defect mesh. Figure 12 (b) shows an energy release event that takes place inside a large amorphous pocket. This event can be viewed as a rearrangement of atoms inside an amorphous pocket. Figure 12 (c) shows an energy release event taking place at a small amorphous zone. This small amorphous zone shrinks even further in this energy release event.
Based on this discovery about the location of annealing events, we can state that the disordered regions in dark matter detector materials, for example, amorphous zones, amorphous-crystalline interface, grain boundaries, interfaces between two different materials, etc., can possibly be active even at cryogenic temperatures. It is likely that defect evolution can take place in these disordered regions, releasing energy that brings disturbance to dark matter direct detection experiments, especially those which are based on phonon detection. This implies that this energy release should be taken into account when building the next generation dark matter detectors. Measures that can be considered to avoid the emergence of atomically disordered regions include optimizing the detector production practice, adopting new manufacturing methods that use less materials with disordered regions, as well as setting up insurmountable barriers between the active detection zone (responsible for interacting with dark matter particles) and other supporting parts of the equipment to block the transmission of the disturbing signals.
VI Estimate of sources and rates of radioactive damage in detectors
As we have shown, the shape of the energy spectrum of annealing events matches well with the excess background observed in the low threshold experiments. In this section we discuss the magnitude of the event rate that could be associated to annealing of radiation damage in silicon detectors. We begin with an estimate of the total power associated with the excess rate, which we obtain by integrating the differential rate shown in Figure 2 of reference [82]. This yields an order of magnitude estimate for the excess power per unit detector mass:
$P_{\rm excess}\approx 10^{5}\frac{\rm eV}{\rm g\,day}.$ (4)
This estimate seems compatible also with the recent observations in two different sized silicon detectors in [21]. A similar excess rate was also observed in a deep underground measurement with a silicon detector in [17].
If the low energy excess is interpreted as arising from the annealing events, the defect concentration must be replenished in the crystal at approximately the same rate as it is being annealed, otherwise the rate would eventually decay. Therefore, the same power $P_{\rm excess}$ needs to be injected and stored in the crystal structure.
There are many known sources of trace radioactivity in dark matter detectors. We consider here some of those that have been observed in experiments and given upper limits [87].
The most common decay modes of the known radioactivity sources are $\alpha$, $\beta$ and $\gamma$ decay. With energies of the emitted particles typically in the MeV range, conservation of momentum implies that also the nucleus from which the recoil occurs receives kinetic energy and may cause damage.
For the $\alpha$ decays, for a known emission energy $E_{kin,\alpha}$ the momentum of the emitted particle is (since the MeV kinetic energies are well below the rest mass of the particles $\gg$ 1 GeV, classical kinematics can be used):
$E_{\rm{kin},\alpha}=\frac{p_{\alpha}^{2}}{2m_{\alpha}}\Rightarrow p_{\alpha}=\sqrt{E_{{\rm{kin}},\alpha}2m_{\alpha}}.$ (5)
a) Refer to caption
b) Refer to caption
c) Refer to caption
Figure 12: The locations of three energy release events. The white meshes are defect meshes identified by Ovito DXA analysis, marking the boundaries of amorphous pockets. By adding colors to the atom displacements in the annealing, the place where atoms move a lot (colored in red) can be easily identified, and it is the movement of atoms in these regions that causes the energy release events. Figure 12(a) is an energy release event taking place at the boundary of an amorphous pocket. Figure 12(b) is an energy release event taking place inside an amorphous pocket. Figure 12(c) is an energy release event taking place at a small amorphous zone. Theta ($\theta$) and phi ($\phi$) give the initial recoil direction in each particular cascade simulation shown here.
Here $p_{\alpha}$ is the momentum and $m_{\alpha}$ the mass of the $\alpha$ particle. Due to momentum conservation, the momentum that the decaying nucleus $A$ receives $p_{A}=p_{\alpha}$ and hence
$E_{{\rm{kin}},A}=\frac{p_{\alpha}^{2}}{2m_{A}}=E_{{\rm{kin}},\alpha}\frac{m_{\alpha}}{m_{A}}.$ (6)
We first consider the decay chain starting from $^{210}{\rm Pb}$ leading to a 5.3 MeV $\alpha$-decay of $^{210}{\rm Po}$. In this process, using eq. 6 one obtains $E_{{\rm{kin}},\rm 206Po}$ = 103 keV [88]. Both the $\alpha$ particle and Pb recoil will in turn lose energy both to electronic and nuclear stopping, out of which only the latter produces significant damage in Si.
To calculate the nuclear deposited energy $F_{D_{n}}$, we used the MDRANGE code [59, 89] in a non-channeling direction using the NLH interatomic potentials [54] to describe the nuclear collisions. We also checked that the SRIM code [90] gave about the same nuclear deposited energies. The MDRANGE code gave $F_{D_{n}}\approx 16$ keV for the $\alpha$ and $F_{D_{n}}\approx 93$ keV for the Po recoil. Hence the total nuclear deposited energy available for damage production is $F_{D_{n}}\approx 110$ keV. Comparison with Figure 4 and the data for Si in Ref. [91] shows that about 8% out of the recoil energy is stored in defects. Thus the stored energy $E_{\rm stored}\approx 9$ keV.
We now use this value to calculate the expected excess power for the possible upper limit of $^{210}{\rm Po}$ activity $A=160\mu{\rm Bq}/{\rm kg}\approx 0.014/{\rm g\,day}$ ([87]). For the Po decays, the predicted excess energy release would be
$P_{\rm excess,210Po}=E_{\rm stored}A\approx 130\frac{\rm eV}{\rm g\,day}.$ (7)
In other words, the Po recoils could explain about 0.1% of the observed excess signal.
We also considered the other parts of the decay chain of 238U. The chain before $^{210}{\rm Pb}$ has in total seven $\alpha$ decays [87], each giving a recoil energy around 100 keV to the parent nucleus. The total recoil energy of these decays is about 670 keV. For the 232Th chain, the five $\alpha$ particles give in total about 490 keV to parent nucleus recoils. The 8% fraction of stored energy $E_{\rm storedU,Th}$ from these recoils is then $\sim 93$ keV. The experimental activity limits for these chains are of the order of $A_{U,Th}\approx 0.001/({\rm g\,day})$ [87, 88], so one can roughly estimate they could contribute
$P_{\rm excessU,Th}=E_{\rm storedU,Th}A_{U,Th}\sim 90\frac{\rm eV}{\rm g\,day},$ (8)
i.e. of the same order as the Po decays, but still not nearly enough to explain the whole signal.
[…]
$P_{\rm excess22Na}=E_{\rm stored22Na}A_{22Na}\sim 0.14\frac{\rm eV}{\rm g\,day}.$ (9)
A similar calculation for the $\beta$ decays of 32Si and tritium 3H gave similarly small values.
At sea level, also cosmic muons and neutrons may produce a lot of damage in the samples; however, at the deep underground conditions where the DM detectors are operated, their direct contribution is negligible [88]. They may indirectly contribute by producing radio-active isotopes such as 32Si when the sample material was on the ground [92]; this contribution is included in the background model [88] that was used to obtain the activities above. We note that a recent observation by the NUCLEUS collaboration [93] finds some apparent difference on the excess rate between surface and underground measurements, but their analysis on the effect does not support the interpretation that the difference could be attributed to cosmic muons.
Finally, we note that in the detectors also decay from components close to the actual detector can be a cause of radioactive background. For instance, for the DAMIC detector, detailed background model calculations showed that events in surrounding copper components dominated the background in some energy intervals and could be comparable to the signal in the detector material and surface (Figure 9 in Ref. [88]). As the goal of the current paper is not to analyze a specific detector, we do not aim at detailed calculations of these effects, but just note that it is reasonable to assume that stored energy caused by radioactivity from the detector surroundings could lead to an excess signal comparable to that from the material itself, i.e. $\sim$ 0.1 - 0.3% of the total.
Taken together, these calculations show that relaxation of energy stored in defects caused by known sources radioactivity in the detectors might explain excess signal powers $\sim$ 200 - 500 eV/(g day) in Si-based detectors, where the lower limit corresponds to our calculation of known sources of radioactivity at the detector material and surface, and the upper limit comes from adding a very rough estimate of the contributions from surrounding materials. This level is experimentally significant, but not sufficient to explain the entire signal observed. Thus it remains to be understood, whether there are other types and sources of damage that could result in similar annealing spectra. For instance, clusters with impurities, inner surfaces of voids, and interfaces between materials are also by definition higher in energy than perfect crystalline bulk, and might be sources of energy release. Systematic consideration of possible energy release from these is beyond the scope of the current study. However, continuing our work, we aim to consider the material-related effects in more detail in the future.
These calculations were all carried out based on the nuclear energy deposition, since in metals and elemental semiconductors, the electronic energy loss does not produce damage. However, we note that in ionic and amorphous materials also the electronic energy loss can lead to major amounts of damage in the form of swift heavy ion tracks [94, 95], and thus in detectors made of ionic materials, similar levels of radioactivity might cause much higher levels of stored energy that may be released.
VII Discussion and conclusion
@@
The main result of this paper is that annealing of damage produced by radioactive events can cause energy release on very long time scales after the damage production, even at cryogenic temperatures. The simulations in this work were done on silicon, but considering that it is well known that other covalently bonded semiconductor materials such as Ge and GaAs produce damage in very similar (and even larger) damage pockets as Si [22, 23, 96, 97], it is very likely that very similar behavior would occur in also other covalently bonded semiconductors. Many ionic materials also undergo amorphization [98, 99, 100], and hence a similar mechanisms might be active in at least some of them. Elemental metals cannot be amorphized [101], however, metastable damage pockets can be formed in them in the form of ordered damage clusters (e.g. C15 structure) and dislocation loops [102, 103, 104, 105]. However, separate simulation works with appropriate interatomic interaction models for other semiconductors, ionic and metallic materials would be needed to deduce whether energy release from other materials would have a similar energy dependence.
@@
The effect described in this paper has a time dependence on the nanosecond to tens of nanoseconds time scale. Most experiments observing the low-energy excess have not reported such time dependence. This is most likely because their time resolution is large (of the order of milliseconds). Hence if there is sufficient amounts of radioactive impurities or any other damage sources in the sample where numerous decays occur in the detector within the time resolution, the average signal on the long time scales would appear constant, because the detector effectively sums up multiple events. On the other hand, we do note that the predicted exponential time dependence of the energy release events on time scales of tens of nanoseconds at room temperature might be experimentally testable with time-resolved pulsed ion beam experiments [106, 107] and sufficiently fast detectors.
Moreover, as we have shown earlier, a dark matter signal could have a distinct daily variation [3, 108, 91], which the excess signals from known sources of radioactive decay would not have. In case dark matter induces damage pockets in materials similar to those studied in the current work, this part would be expected to have exactly the same energy release behavior as the ones studied here. Hence, while regular matter effects appear to dominate the low energy excess signal, some fraction of it might in principle also have a dark matter origin, and this part would also be expected to have a daily variation.
@@
In conclusion, using molecular dynamics simulations of radiation events and their annealing in silicon, we have shown that even at cryogenic temperatures, energy release events of up to 10’s of electron volts can be caused by random annealing of complex disordered atom pockets. We show that the annealing can occur by avalanche-like events, where a metastable defect configuration first crosses an energy barrier of about 0.1 eV. Crossing this kind of small barrier can then trigger a rapid additional bond rearrangements, that lead to partial recrystallization and energy releases exceeding 10 eV. This explains why large energy release events can occur even down to cryogenic temperatures. These events occur randomly up to macrosopic time scales after the initial picosecond time scale radiation event. Hence, in a sensitive detector they could be caused by annealing of damage produced e.g. by decay of radioactive impurities either in the detector itself or in materials surrounding it. These energy release events have an exponential energy dependence that agrees very well with the experimentally observed low-energy excess energy release in semiconductor detectors. Estimation of known sources of radioactivity in high-quality Si detectors showed that their concentration can explain at most $\sim$ 0.5% of the observed excess signal, showing that further work is needed to find whether the remaining signal is caused by other structures storing energy, or is due to some completely different effects that also happen to have a similar energy release spectrum.
@@
Appendix A Verifying the settings of quantum thermal bath
There are five parameters to set the quantum thermal bath, which are the target quantum temperature, the temperature damping parameter, the seed for the random number generator to create the random forces, the cutoff frequency of the vibration spectrum, and the number of frequency bins used to discretize the frequency range.
The target quantum temperature is the required temperature for the system, being set at the beginning of annealing simulations. The temperature damping parameter is the reciprocal of the friction coefficient $\gamma$, deciding how rapidly the temperature is relaxed to the target temperature. According to LAMMPS manual [109], the suggested range of temperature damping parameter is 0.2 ps to 1 ps, making the friction coefficient $\gamma$ fall into the range of 1 THz to 5 THz. The seed for the random number generator is a positive integer, being used to create the random forces. The cutoff frequency determines the upper limit of the frequency domain of the random forces, and vibrational modes with higher frequencies will not be simulated. According to LAMMPS manual [109], the recommended range of cutoff frequency is 2 to 3 times of the Debye frequency of the simulated system. The number of frequency bins sets how many points are sampled in the frequency domain. According to LAMMPS manual [109], the number of frequency bins should be large enough so that the classical temperature can converge to the quantum temperature as the temperature increases.
Among these five parameters, the temperature damping parameter and the cutoff frequency have physical significance, so these two parameters should be tested, evaluating how much influence they have on calculation results.
First, the influence of temperature damping is investigated. Figure 13 demonstrates the heat capacity calculation results with different temperature damping parameters. The three heat capacity curves from different temperature damping parameters overlap each other. One can see from this figure that as long as the temperature damping falls in the suggested range given by LAMMPS manual, the influence of temperature damping on the heat capacity results is minimal. Also, all these three heat capacity curves agree with the experimental result (given by Anderson et al. [67]) quite well.
Figure 14 shows the relationship between the classical temperature and the quantum temperature, demonstrating the influence of temperature damping on it. In this figure, the three temperature curves are very close to each other. One can see that as long as the suggestions from LAMMPS manual is followed, the temperature curve is not sensitive to temperature damping parameter. Therefore, the temperature damping chosen in the simulations with quantum thermal bath in this paper is 0.3 ps, being identical with the temperature damping used in classical thermostats in this paper, enabling a direct comparison between quantum thermal bath results and classical thermostat results.
Figure 13: The influence of temperature damping parameter (“damp” in the figure) on heat capacity results
Figure 14: The influence of temperature damping parameter (“damp” in the figure) on the relationship between quantum temperature and classical temperature
Second, the influence of cutoff frequency is evaluated. Figure 15 demonstrates the heat capacity calculation results with different cutoff frequencies. The three heat capacity curves from different cutoff frequencies overlap each other. This means that as long as the requirements in LAMMPS manual are followed, the heat capacity results are not sensitive to the cutoff frequency. Also, all of the three curves agree with the experimental result from Anderson et al. [67] quite well.
Figure 16 shows the relationship between the classical temperature and the quantum temperature, demonstrating the influence of cutoff frequency on it. In this figure, the three temperature curves from different cutoff frequencies are very close to each other. As long as one follows the suggestions from LAMMPS manual, the temperature curve is not sensitive to cutoff frequency. Therefore, the cutoff frequency chosen in the simulations with quantum thermal bath in this paper is 3 times the Debye frequency of silicon. The reason for choosing this value is that Stillinger-Weber potential overestimates the Debye frequency of silicon by around 17% [66]. The input cutoff frequency is calculated by three times the Debye frequency coming from the Debye temperature of silicon, which is from reference [83]. This value is smaller than three times the Debye frequency of Stillinger-Weber potential, falling in the suggested range from LAMMPS manual, 2 to 3 times the Debye frequency of the system (Stillinger-Weber potential).
Figure 15: The influence of cutoff frequency ($f_{max}$ in the figure) on heat capacity results
Figure 16: The influence of cutoff frequency ($f_{max}$ in the figure) on the relationship between quantum temperature and classical temperature
Then, with the temperature damping being 0.3 ps and cutoff frequency being 3 times the Debye frequency of silicon, the relationship between heat capacity and temperature is calculated, which is shown in Figure 17. From this figure, the heat capacity curve calculated with quantum thermal bath agrees well with experimental results given by Anderson et al. [67] at low temperatures. Moreover, the heat capacity curve converges to the classical limit given by Dulong-Petit Law at high temperatures. Therefore, one can see that with these settings, the temperature dependence of heat capacity can be correctly reproduced.
Figure 17: The heat capacity calculation result when the temperature damping parameter is 0.3 ps and the cutoff frequency is 3 times the Debye frequency of silicon.
Finally, the relationship between the classical temperature and the quantum temperature is calculated at a large temperature range from 0 K to 800 K, with the temperature damping being 0.3 ps and cutoff frequency being 3 times the Debye frequency of silicon. The result is shown in Figure 18.
Figure 18: The relationship between quantum temperature and classical temperature when the temperature damping parameter is 0.3 ps and the cutoff frequency is 3 times the Debye frequency of silicon. The number of frequency bins is 100.
In this figure, a non-zero classical temperature is present even when the quantum temperature approaches zero. This is associated with zero-point energy at low temperatures. At high temperature, the classical temperature gradually converges to the quantum temperature, which is the classical limit. According to LAMMPS manual [109], this convergence also justifies the setting of the number of frequency bins ($N_{f}$). In the simulations with quantum thermal bath in this paper, the number of frequency bins is set to be 100. This convergence in Figure 18 means that $N_{f}=100$ is large enough to obtain the convergence of the classical temperature to the target quantum temperature in a high temperature regime.